OpenStudy (anonymous):

Determine where, if anywhere, the tangent line to f(x) = x^3 - 5x^2 + x is parallel to the line y = 4x +23.

OpenStudy (phi):

can you find the derivative of f(x)? if so, set the derivative equal to 4 (the slope of the line) and solve for x

OpenStudy (anonymous):

if we know that the first derivative of f(x) is the slope of the tangent, then all we need to do is set this Fprime equl to 4

OpenStudy (anonymous):

@phi Why set the derivative equal to 4?

OpenStudy (phi):

you want a line parallel to the line y = 4x +23. parallel means equal slopes. 4 is the slope of the given line (y = mx+b format) as you may know, the derivative of a curve will give you the slope of the tangent line to the curve at any point.

OpenStudy (anonymous):

@phi Then would I just simply simplify the equation and that will be the line?

OpenStudy (phi):

after the derivative, you get a quadratic. solve for x. if you get real numbers, then that is the answer. if you get complex numbers, there is no tangent that is parallel to the given line.

OpenStudy (anonymous):

I'm confused on the quadratic part of the equation.

OpenStudy (phi):

did you get the derivative?

OpenStudy (anonymous):

Yes, I got 3x^2 - 10x + 1

OpenStudy (phi):

so set = to 4: 3x^2 - 10x + 1=4 or 3x^2 - 10x -3= 0 use the quadratic formula to solve. http://www.purplemath.com/modules/quadform.htm

OpenStudy (anonymous):

I have \[(10 \pm \sqrt{136} ) \div 6\]

OpenStudy (phi):

so you have 2 x values where the tangent line will be parallel to the given line. You can find the corresponding y value (use the original equation) to give 2 (x,y) pairs as the answer to the question.

OpenStudy (anonymous):

@phi I understand now! Thank you very much!

OpenStudy (anonymous):

@phi But would I use the f(x) equation or the y equation?

OpenStudy (phi):

the f(x) equation (you are looking for a tangent to f(x), so the (x,y) point has to be on that curve) the other line y= 4x+23 is parallel to the tangent , but not the same as the tangent line.

OpenStudy (anonymous):

@phi Thank you!